# On the Effectiveness of Membership Inference in Targeted Data Extraction from Large Language Models

### v1

@@ II-B Membership Inference Attacks Specifically for LLMs
Offering a more nuanced perspective, Chen et al. [27] provided a statistical and multi-perspective revisiting of MIAs, finding that while overall performance is often low, a significant number of ”differentiable outliers” exist, where member and non-member instances are easily distinguishable . Their work suggests that the inconsistency in prior research stems from the high dependency of MIA success on specific contexts like data domain and model size, rather than a universal failure of the attacks. Hayes et al. [28] scaled one of the strongest MIAs (LiRA) to models up to 1B parameters. Their findings confirmed that strong attacks can succeed, but their effectiveness is limited in practical, single-epoch training scenarios. They also demonstrated that a sample’s vulnerability to membership inference does not directly correlate with its risk of extraction, suggesting these two privacy threats capture different aspects of memorization.
@@ III-A Defining Memorization
Definition III.1 ($k$-eidetic memorization [18])
@@ III-B Basic Setups
Following the setup of the Data Extraction Challenge, we employ the GPT-Neo-1.3B model [33], an open-source transformer model designed by training EleutherAI’s replication of the GPT-3 architecture on The Pile dataset. GPT-Neo is an autoregressive language model $f_{\theta}$ that generates a sequence of tokens $x_{0},x_{1},\ldots,x_{N}$ via the chain rule:
@@ III-C Threat Model
We consider an adversary with black-box access to the target LLM. In this setting, the adversary can query the model with arbitrary inputs (prompts or prefixes) and observe the generated text outputs along with the associated probabilities or log-likelihoods of the tokens. However, the adversary cannot access the model’s internal parameters (weights), architecture details beyond publicly available information, or the training dataset.
@@ III-D Evaluation Metrics
$\begin{split}M_{P}=\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}[\hat{y}_{i}^{(1)}=y_{i}],\end{split}$ (3)
@@
$\begin{split}M_{H}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{1}{L}\sum_{j=1}^{L}\mathbbm{1}[x_{i,j}\neq y_{i,j}]\right),\end{split}$ (4)
where $x_{i,j}$ and $y_{i,j}$ denote the $j$-th tokens of the generated ($x_{i}=\hat{y}_{i}^{(1)}$) and ground-truth ($y_{i}$) suffixes of length $L$, respectively.
@@ IV-C Ranking Methods
$\begin{split}M_{\text{Zlib}}(s|p)=\frac{\text{LL}(s|p)}{\text{Zlib}(s)}\end{split}$ (6)
@@
$\begin{split}M_{\text{HC}}(s|p)=\frac{1}{|s|}\sum_{x_{t}\in s}\mathcal{L}_{\text{adj}}(x_{t})\end{split}$ (7)
@@
Outlier-Robust Likelihood [31]: To prevent single outlier tokens from disproportionately influencing the score, Yu et al. introduced this approach which calculates Likelihood using only raw tokens whose log-probabilities fall within a specified range, replacing outliers with the mean of all tokens in the sequence.
@@
$M_{\text{ORL}}(s|p)=\frac{1}{|s|}\sum_{x_{t}\in s}\mathcal{L}^{\prime}(x_{t})$ (10)
SURP (Surprising Tokens) [40]: Zhang et al. introduced this method, which identifies ”surprising tokens” in $s$ and uses their average log-probability as the membership score. A token is considered surprising if the model is confident in its prediction, but assigns a low probability to the actual token. The score is higher for sequences where these surprising tokens are less surprising. Let $\mathcal{T}_{\text{SURP}}$ be the set of surprising tokens in $s$.
$\begin{split}M_{\text{SURP}}(s|p)=\frac{1}{|\mathcal{T}_{\text{SURP}}|}\sum_{x_{t}\in\mathcal{T}_{\text{SURP}}}\log P(x_{t}|p,x_{<t})\end{split}$ (11)
ReCaLL (Relative Conditional Log-Likelihood) [24]: Xie et al. found that conditioning the target on a generic non-member prefix causes a noticeable log-likelihood shift which can be exploited for membership inference. This method examines the ratio of the conditional log-likelihood of the suffix $s$ when prefixed by a non-member context $p_{nm}$ to its unconditional log-likelihood.
$\begin{split}M_{\text{R}}((p,s);p_{nm})=\frac{\text{LL}((p,s)|p_{nm})}{\text{LL}((p,s))}\end{split}$ (12)
Suffix ReCaLL (S-ReCaLL): Built on the premise that a true prefix should be highly predictive of its memorized suffix, we calculate the ratio of the suffix’s unconditional negative log-likelihood (NLL) to its conditional NLL given the true prefix $p$. This represents an extension of the ReCaLL method designed to exploit the true prefix as an additional source of information.
$\begin{split}M_{\text{SR}}(s|p)=\frac{\text{LL}(s|p)}{\text{LL}(s)}\end{split}$ (13)
@@
$\begin{split}M_{\text{CR}}(s|p;p_{nm})=\frac{\gamma\cdot\text{LL}((p,s)|p_{m})-\text{LL}((p,s)|p_{nm})}{\text{LL}((p,s))}\end{split}$ (14)
Lowercase [18]: Carlini et al. introduced this metric which measures the ratio of a model’s perplexity on an original suffix compared to its fully lowercased version. The intuition is that verbatim content often preserves distinctive capitalization patterns, and when these are normalized, the model assigns them a substantially lower likelihood, thereby signaling memorization.
Min-k% Prob (Min-K%) [20]: Shi et al. based this method on the claim that even the least predictable parts of a memorized sequence should have high likelihood. The method calculates the average log probability of the $k\%$ tokens in $s$ (given $p$) with the lowest conditional probabilities. A higher score indicates stronger confidence even in the ”hardest” parts of the sequence.
$\begin{split}M_{\text{K\%}}(s|p)&=\frac{1}{|\mathcal{T}_{\min\text{-}k\%}|}\\ &\quad\times\sum_{x_{t}\in\mathcal{T}_{\min\text{-}k\%}}\log P(x_{t}|p,x_{<t})\end{split}$ (15)
@@
$\begin{split}M_{\text{K\%+}}(s|p)&=\frac{1}{|\mathcal{T}_{\min\text{-}k\%}|}\\ &\quad\times\sum_{x_{t}\in\mathcal{T}_{\min\text{-}k\%}}\left(\frac{\log P(x_{t}|p,x_{<t})-\mu_{p,x_{<t}}}{\sigma_{p,x_{<t}}}\right)\end{split}$ (16)
@@ V-C Tuning Hyperparameters for MIAs
Figure 3: Performance of various MIA methods as their hyperparameters are tuned in a standard MIA setting. (a) shows probability-based methods as threshold ’k’ changes. (b) shows ReCaLL-based methods as the number of non-member samples ’N’ changes.
@@ V-D Results on the Extended Subset
Unlike the ranking stage where most methods performed almost identically, here we see a more distinct separation. While S-ReCaLL is the clear top performer, several other methods, including Min-K% (AUROC 84.6%) and the simple Likelihood baseline (AUROC 83.3%), prove to be surprisingly competitive. This indicates that raw model confidence remains a strong signal for this classification task. Methods like Lowercase, ReCaLL, and especially Min-K%++ perform poorly, confirming their unsuitability for this confirmation step.
@@ VI Targeted Extraction from Fine-Tuned Models
While the previous sections analyzed extraction from pre-trained base models, a fundamental limitation in evaluating privacy risks on such models is the lack of transparency regarding their training data. For most state-of-the-art LLMs, the training corpora are not open-source, making it difficult for researchers to establish the ”ground truth” of memorization—specifically. By performing controlled fine-tuning experiments, we can precisely control the injection of sensitive data and evaluate the attack against a wider range of model architectures.
@@ VI-B Impact of Repetition on Memorization
We first evaluate the extraction success rate, which is the percentage of phone numbers correctly recovered by the model given the prefix. As shown in Table VIII, even a single occurrence in the fine-tuning data creates a substantial privacy risk.
@@
TABLE VIII: Extraction Success Rate (Phone Number Match %) by Repetition Count
@@ VII Discussion
This paper is organized around two main parts of the data extraction pipeline. The first part provides a systematic benchmark of MIA techniques as ranking functions within the standard two-stage pipeline. The second part evaluates the use of MIA-based filtering to prune false positives from the final output.
@@
A similar trend is observed in the MIA-based confirmation stage: most methods yield only marginal gains over the baseline likelihood, unlike the gains achieved in MIA benchmarks.

### v2

@@ II-B Membership Inference Attacks Specifically for LLMs
Offering a more nuanced perspective, Chen et al. [27] provided a statistical and multi-perspective revisiting of MIAs, finding that while overall performance is often low, a significant number of "differentiable outliers" exist, where member and non-member instances are easily distinguishable. Their work suggests that the inconsistency in prior research stems from the high dependency of MIA success on specific contexts like data domain and model size, rather than a universal failure of the attacks. Hayes et al. [28] scaled one of the strongest MIAs (LiRA) to models up to 1B parameters. Their findings confirmed that strong attacks can succeed, but their effectiveness is limited in practical, single-epoch training scenarios. They also demonstrated that a sample’s vulnerability to membership inference does not directly correlate with its risk of extraction, suggesting these two privacy threats capture different aspects of memorization.
@@ III-A Defining Memorization
Definition 1 ($k$-eidetic memorization [18])
@@ III-B Basic Setups
Following the setup of the Data Extraction Challenge, we employ the GPT-Neo-1.3B model [33], An open-source transformer replicating GPT-3, trained by EleutherAI on The Pile. GPT-Neo is an autoregressive language model $f_{\theta}$ that generates a sequence of tokens $x_{0},x_{1},\ldots,x_{N}$ via the chain rule:
@@ III-C Threat Model
We consider an adversary with black-box access to the target LLM. In this setting, the adversary can query the model with arbitrary inputs (prompts or prefixes) and observe the generated text outputs along with the associated probabilities or log-likelihoods of the tokens. However, the adversary cannot access the model’s weights, architecture details beyond publicly available information, or the training dataset.
@@ III-D Evaluation Metrics
$\begin{split}M_{P}=\frac{1}{N}\sum_{i=1}^{N}\mathds{1}[\hat{y}_{i}^{(1)}=y_{i}],\end{split}$ (3)
@@
$\begin{split}M_{H}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{1}{L}\sum_{j=1}^{L}\mathds{1}[x_{i,j}\neq y_{i,j}]\right),\end{split}$ (4)
where $x_{i,j}$ and $y_{i,j}$ denote the $j$-th tokens of the top-1 generated suffix ($x_{i}=\hat{y}_{i}^{(1)}$) and ground-truth suffix ($y_{i}$) of length $L$, respectively.
@@ IV-C Ranking Methods
$\begin{split}M_{\text{Zlib}}(s|p)=\frac{\text{LL}(s|p)}{\text{Zlib}(s)}.\end{split}$ (6)
@@
$\begin{split}M_{\text{HC}}(s|p)=\frac{1}{|s|}\sum_{x_{t}\in s}\mathcal{L}_{\text{adj}}(x_{t}),\end{split}$ (7)
@@
Outlier-Robust Likelihood [31]: To prevent outlier tokens from disproportionately influencing the score, Yu et al. introduced this approach which calculates Likelihood using only raw tokens with log-probabilities that fall within a specified range, and replaces remaining outliers with the mean of all tokens in the sequence.
@@
$M_{\text{ORL}}(s|p)=\frac{1}{|s|}\sum_{x_{t}\in s}\mathcal{L}^{\prime}(x_{t}).$ (10)
SURP (Surprising Tokens) [40]: Zhang et al. introduced this method, which identifies "surprising tokens" in $s$ and uses their average log-probability as the membership score. A token is considered surprising if the model is confident in its prediction, but assigns a low probability to the actual token. The score is higher for sequences where these surprising tokens are less surprising. Let $\mathcal{T}_{\text{SURP}}$ be the set of surprising tokens in $s$.
$\begin{split}M_{\text{SURP}}(s|p)=\frac{1}{|\mathcal{T}_{\text{SURP}}|}\sum_{x_{t}\in\mathcal{T}_{\text{SURP}}}\log P(x_{t}|p,x_{<t}).\end{split}$ (11)
ReCaLL (Relative Conditional Log-Likelihood) [24]: Xie et al. found that conditioning the target on a generic non-member prefix causes a noticeable log-likelihood shift. This method examines the ratio of the conditional log-likelihood of the suffix $s$ when prefixed by a non-member context $p_{nm}$ to its unconditional log-likelihood.
$\begin{split}M_{\text{R}}((p,s);p_{nm})=\frac{\text{LL}((p,s)|p_{nm})}{\text{LL}((p,s))}.\end{split}$ (12)
S-ReCaLL( Suffix ReCaLL ): Built on the premise that a true prefix should be highly predictive of its memorized suffix, we introduce this method which calculates the ratio of the suffix’s unconditional negative log-likelihood (NLL) to its conditional NLL given the true prefix $p$. This represents an extension of the ReCaLL method designed to exploit the true prefix as an additional source of information.
$\begin{split}M_{\text{SR}}(s|p)=\frac{\text{LL}(s|p)}{\text{LL}(s)}.\end{split}$ (13)
@@
$\begin{split}M_{\text{CR}}(s|p;p_{nm})=\frac{\gamma\cdot\text{LL}((p,s)|p_{m})-\text{LL}((p,s)|p_{nm})}{\text{LL}((p,s))}.\end{split}$ (14)
Lowercase [18]: Carlini et al. introduced this metric which measures the ratio of a model’s perplexity on an original suffix compared to its fully lowercased version. The intuition is that verbatim content often preserves distinctive capitalization patterns, and when these are normalized, the model is expected to assign them a substantially lower likelihood, thereby signaling memorization.
Min-k% Prob (Min-K%) [20]: Shi et al. based this method on the claim that even the least predictable parts of a memorized sequence should have high likelihood. The method calculates the average log probability of the $k\%$ tokens in $s$ (given $p$) with the lowest conditional probabilities. A higher score indicates stronger confidence even in the "hardest" parts of the sequence.
$\begin{split}M_{\text{K\%}}(s|p)&=\frac{1}{|\mathcal{T}_{\min\text{-}k\%}|}\\ &\quad\times\sum_{x_{t}\in\mathcal{T}_{\min\text{-}k\%}}\log P(x_{t}|p,x_{<t}).\end{split}$ (15)
@@
$\begin{split}M_{\text{K\%+}}(s|p)&=\frac{1}{|\mathcal{T}_{\min\text{-}k\%}|}\\ &\quad\times\sum_{x_{t}\in\mathcal{T}_{\min\text{-}k\%}}\left(\frac{\log P(x_{t}|p,x_{<t})-\mu_{p,x_{<t}}}{\sigma_{p,x_{<t}}}\right).\end{split}$ (16)
@@ V-C Tuning Hyperparameters for MIAs
Figure 3: Performance of various MIA methods as their hyperparameters are tuned in a standard MIA setting.
@@ V-D Results on the Extended Subset
Unlike the ranking stage where most methods performed almost identically, here we see a more distinct separation. While S-ReCaLL is the top performer, several other methods, including Min-K% (AUROC 84.6%) and Likelihood (AUROC 83.3%), remain competitive. This indicates that raw model confidence remains a strong signal for this classification task. Methods like Lowercase, ReCaLL, and especially Min-K%++ perform poorly, confirming their unsuitability for this confirmation step.
@@ VI Targeted Extraction from Fine-Tuned Models
While the previous sections analyzed extraction from pre-trained base models, a fundamental limitation in evaluating privacy risks on such models is the lack of transparency regarding their training data. For most state-of-the-art LLMs, the training corpora are not open-source, making it difficult for researchers to establish the "ground truth" of memorization—specifically. By performing controlled fine-tuning experiments, we can precisely control the injection of sensitive data and evaluate the attack against a wider range of model architectures.
@@ VI-B Impact of Repetition on Memorization
We first evaluate the extraction success rate, which is the percentage of phone numbers correctly recovered by the model given the prefix. As shown in Table VIII, even a single occurrence in the fine-tuning data poses a significant privacy risk.
@@
TABLE VIII: Extraction Success Rate by Repetition Count
@@ VII Discussion
This paper is organized around two main parts of the data extraction pipeline. The first part provides a systematic benchmark of MIA techniques as ranking functions within the standard two-stage pipeline. The second part evaluates the use of MIA-based filtering to prune false positives from the final extracted output.
@@
A similar trend is observed in the MIA-based confirmation stage: most methods yield only marginal gains over the baseline likelihood, unlike the gains achieved in MIA benchmarks used in related work.